In 28 years Mark will be 5 times older that he is right now. What is Mark's current age?

Answer by checkley79(3341)

X+28=5X
X-5X=-28
-4X=-28
X=-28/-4
X=7 YEARS OLD NOW FOR MARK.
7+28=5*7
35=35